OpenStudy (anonymous):

Evaluate: anti-deriviative 3/(x-4) dx will I use U substitute

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Identify your composite functions f(a) and g(x) such that 3/(x-4) = f(g(x))

OpenStudy (anonymous):

u= x-4 du= 1

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Well, actually not quite

OpenStudy (anonymous):

du = 1dx

OpenStudy (anonymous):

yes I had that just forgot to type it

OpenStudy (anonymous):

ok good.

OpenStudy (anonymous):

ok this is where I have trouble.....will it be 3/u ???? Doesn't look right

OpenStudy (anonymous):

That is exactly right. If u = x-4 then 3/(x-4) = 3/u

OpenStudy (anonymous):

so will I add 1 to the u (exponent) to be u^2

OpenStudy (anonymous):

Not quite. realize that you basically have \[3 * \frac{1}{u}\] So the derivative of what function of u gives you \(\frac{1}{u}\)?

OpenStudy (anonymous):

u^-1 ??

OpenStudy (anonymous):

1/u is u^-1

OpenStudy (anonymous):

So it's not that. \(\frac{d}{du}[ln\ u] = \frac{1}{u}\) remember?

OpenStudy (anonymous):

not sure what you mean about that?

OpenStudy (anonymous):

The derivative of ln x = 1/x right?

OpenStudy (anonymous):

ln is the natural logarithm function.

OpenStudy (anonymous):

oh, yes it does. Too many rules.....

OpenStudy (anonymous):

yeah, you'll get them with practice.

OpenStudy (anonymous):

So the anti derivative of 1/x = ln x + C

OpenStudy (anonymous):

So 3 times the anti derivative of 1/u du = ?

OpenStudy (anonymous):

ok lost again..... I know we have 3/u or 3 * 1/u would it be 3lnx + c ???

OpenStudy (anonymous):

It would be 3(ln u) + C, but then you have to plug in what u is.

OpenStudy (anonymous):

how about 3 ln\[\left| x-4 \right|\] + C sorry couln't get them all in one line

OpenStudy (anonymous):

Yep, that's it exactly.

OpenStudy (anonymous):

I am getting there......slowly....... Thanks I really appreciate your help.

OpenStudy (anonymous):

Of course! Just keep practicing. You're doing very well

OpenStudy (anonymous):

I have a test next Tue. So I will be asking questions up until then. I am sure I will talk to you again..

OpenStudy (anonymous):

ok!

OpenStudy (anonymous):

I got really good with the derivatives, so my brain still wants to do them.

OpenStudy (anonymous):

Yeah, I know the feeling. Soon you'll get the hang of going the other way.

OpenStudy (anonymous):

I will start one more question before I have to leave

OpenStudy (anonymous):

ok